OpenStudy (anonymous):

At Priscilla's Purse Warehouse, it takes 2/3 of a day to complete 1/9 of an order of purses. At this rate, how long will it take to complete the entire order of purses?

hero (hero):

Hint: (1/9)*9 = 1 order of purses

OpenStudy (anonymous):

thx

hero (hero):

@Maxipoo124, did you figure how how long it will take to complete the entire order?

OpenStudy (anonymous):

yes

hero (hero):

How long will it take?

OpenStudy (anonymous):

13 1/2 days

hero (hero):

What calculations did you use to come up with that result?

OpenStudy (anonymous):

2/3 x 9 cross multiplied

hero (hero):

There's no cross multiplying involved in calculating (2/3)*9 ,

OpenStudy (anonymous):

well i multiplied 3 x 9 and divided it by 2 and got 13.5 but into a fraction its 13 1/2

hero (hero):

You only cross multiply if you have a proportion: \[\dfrac{a}{b} = \dfrac{c}{d}\] \[ad = bc\]

hero (hero):

This is what you're multiplying: \[\dfrac{2}{3} \times \dfrac{9}{1}\] You multiply the numerators first. Then the denominators

OpenStudy (anonymous):

so six days

hero (hero):

Yes, correct.

OpenStudy (anonymous):

thank you so much for the help